Beth C. Day lived 63 years from 18tu to 18ut. The sum of tu and ut is 11. In what year was Beth C. Day born?
explain why
18tu-18ut=63; to
1800+10t+u-(1800+10u+t)=63; - why 10?

Answer by richwmiller(17219)

t=2 u=9
1892-1829=63
(1800+10u+t)-(1800+10t+u)=63
why 10?
29 is 10*2+9
92 is 10*9+2